The average grade in a statistics course has been 76 with a standard deviation of 11. If a random sample of 56 is selected from this population, what is the probability that the average grade is more than 80? Use Appendix B.1 for the z-values. (Round your z-value to 2 decimal places and the final answer to 4 decimal places.)

Probability           
Answer #1

Solution :

Given that ,

mean = \mu = 76

standard deviation = \sigma = 11

\sigma\bar x = \sigma / \sqrt n = 11 / \sqrt 56 = 1.4699

P(\bar x > 80) = 1 - P(\bar x < )80

= 1 - P[(\bar x - \mu \bar x ) / \sigma \bar x < (80 - 76) / 1.4699]

= 1 - P(z < 2.72)

= 1 - 0.9967

= 0.0033

Probability = 0.0033